# THERADIA WoZ: An Ecological Corpus for Appraisal-based Affect Research in Healthcare

### v1

@@ VI-A1 Expert based representations
We used the Mel-scale Filter Banks (MFBs), Term Frequency - Inverse Document Frequency (TF-IDF), and Facial Action Units (FAUs), as representations of the audio, textual, and visual data, respectively.
@@ VI-A2 Self-supervised representations
We used a multilingual Wav2Vec2 model that exploits both convolutional and transformer layers to distinguish positive and negative samples of speech in a contrastive approach [46]. The model [^7] https://huggingface.co/voidful/wav2vec2-xlsr-multilingual-56 we used has a large architecture, and was trained on 56 different languages [53], including French, and then fine-tuned for ASR.
@@ VI-B Models and training strategies
In order to predict the time-continuous annotations of the dimensions, we had to re-sample some features to match the annotation frame rate, which was 10 Hz. Audio representations, which had a sampling rate of 100 Hz for MFBs, and 50 Hz for Wav2vec2, were simply averaged into chunks of 10 ms to match the sampling frequency of the annotations; the same strategy was applied to the FAUs and CLIP visual representations. Regarding textual representations, we only exploited BERT as the TF-IDF returns a single feature vector for the whole sequence. Textual features were aligned to the size of the sequence using a spline interpolation, which provided a smooth and continuous representation of the features over time. As the Linear and MLP models did not achieve good performance for the prediction of time-continuous dimensions, we further experimented with various GRU models:
@@
    One-Layer GRU: A single-layered GRU with the hidden size equal to half of the input feature size, or equal to the input feature size.
@@
    Three-Layers GRU: Similar to before, but with three hidden layers. A set of fixed hidden sizes was also tested: 128, 256, and 512.
Results showed that a GRU model with three hidden layers of the same size as the input was optimal for deep representations, whereas the same model with a hidden size of 256 yielded the best results for the hand-crafted representations.
For multimodal fusion, we experimented with middle and late fusion techniques to deal with the asynchronous inputs. The best performing approach for both labels and dimension summaries was a decision-based fusion using an Ordinary Least Squares (OLS) regression model [^10] https://scikit-learn.org. Fusion of the time-continuous dimensions was operated with a GRU as a weighted averaging layer to take into account the temporal dynamics of the features. The weights of the GRU were optimised on the training set.
All the models were trained with the Adam optimizer and an initial learning rate of $10^{-3}$ for affective labels and $10^{-4}$ for dimensions. The batch size and gradient accumulation were set to 10 for labels and 1 for dimensions. The maximum number of training epochs was set to 50 and an early stopping strategy was used with five epochs.Training of the models was done using the Mean Squared Error (MSE) as loss function, and for evaluation, the Concordance Correlation Coefficient (CCC) was calculated on all the sequences [56]. All experiments were carried out using Pytorch [^11] https://pytorch.org[57], with random seeds set to zero. The computer’s operating system was Debian GNU/Linux 10, and the GPU used to train the models was an NVIDIA Quadro RTX 6000 with 23 gigabytes of memory, CUDA version 11.3.
@@ VI-C Affect prediction for the core set labels
Results obtained on the intensity of the labels are reported in Table V and show some strong dependencies between the modalities and the affective labels (see section III.B of the supplementary material for further details on prediction bias based on population demographics). For example, the label “happy” was more accurately evaluated by the video modality, the label “interested” by the textual modality, and the labels “frustrated” and “surprised” by the audio modality with deep representations. The choice of using either expert or deep features may further lead to important differences in the results according to the modality. For example, “frustrated” was detected with a much higher CCC (.327) with the TF-IDF compared to MFB (.151), but the opposite is observed when using the SSL representations. On average, deep representations of the textual, audio, and video modalities performed significantly better compared to their hand-crafted representations.
The fusion of the different modalities provided the best results on average, and for most affective labels. An analysis of the fusion weights showed that the relevance of each modality strongly depends on its representation. For example, when using expert features, the TF-IDF presents a much greater influence on the results compared to the MFBs and FAUs, with an averaged fusion coefficient of 0.8, compared to 0.1 for audio and video, whereas the contribution of those two modalities outweigh that of text when using deep representations, with averaged coefficients of .28, .35, and .35 for text, audio, and video respectively.
TABLE V: Core-set label intensity prediction results (CCC).
@@
An affective label was defined as present for a given sequence if at least one annotator had reported an intensity value for this label. We investigated two modeling options for their prediction: learning a binary classification model for each affective label, or learning a threshold from the regression models. As the binary classification task achieved significantly worse results compared to a decision based on the regression models, we therefore chose the latter for the experiments.
Results are shown in table VI. Similar to the previous experiments, there exists some dependencies between the best performing modality and the affective labels, even though the difference is less contrasted. Overall, we do not observe a significant difference between the expert and deep representations for the different modalities, excepted for the audio modality, and the fusion improved the performance almost systematically.
@@ VI-D Dimensions summaries prediction
TABLE VI: Core-set label classification results (%UAR).
@@
Results are reported in table VII for two different gold standards: the average of annotations and the median (see section III.B of the supplementary material for further details on prediction bias based on population demographics). Comparing those two gold standards, we notice that the average of the annotations yielded the best performance across the five dimensions. Regarding the dimensions, arousal was best predicted by the audio modality and the two others modalities did not improve the performance during fusion. Arousal also achieved the lowest scores compared to the others dimensions, which might be explained by the reasons given in V-C. Coping was the best predicted dimension where all modalities performed similarly well and showed a complementarity during fusion. Goal conduciveness was best predicted by the textual representations, and fusion with the audio and video modalities also improved performance. Intrinsic pleasantness was best captured by the video modality; textual representations were relevant only when computed at the whole level of the sequence with the TF-IDF, and deep representations of audio performed similarly well. Novelty appeared to be really challenging for the video modality with extremely low scores, whereas textual representations captured better this dimension.
By analysing the relative importance of each modality based on the coefficients obtained from the linear regression model, we observed that a very low importance was given to audio features for arousal, in particular when using expert features. This may explain the decrease in performance for the multimodal approach, compared to using the audio modality alone. For the others dimensions, it was observed that the contributions of the modalities were influenced by the complexity of the representations. When using deep features, the audio modality emerged as the most influential, but this contribution was reduced for expert features, with textual information showing more important contributions.
TABLE VII: Dimension summary prediction (CCC).
@@ VI-E Continuous dimensions prediction
Results are reported in table VIII for three gold standards: the average and the median of the time-continous annotations, and a version based on the Evaluator Weighted Estimator (EWE), which corresponds to a weighted mean of the annotations using the average of the cross-CCC between each rater’s as weights [58] (see section III.B of the supplementary material for further details on prediction bias based on population demographics). Overall, the EWE approach systematically outperformed the others two gold standards on the five dimensions.
Arousal was the best predicted dimension, but this is due to the low variance observed within this dimension, as explained in the previous section. For coping, goal conduciveness and novelty, the Wav2vec2 audio representations performed better than the textual features, differently from what was observed for the summarised dimensions, which might be due to the fact that textual features provide a static representation that may not capture the continuous variations within the data. Regarding intrinsic pleasantness, the three modalities performed similarly and proved complementary during fusion, which is consistent with the observations made in the previous section. In addition, the multimodal approach consistently improved the performance. An analysis of the fusion weights revealed that when using the mean of annotations as the gold standard, the model evenly distributed importance across all modalities. However, when using the median or the EWE gold standards, the audio modality appeared as the most significant, whereas text and video contributed similarly.
An examination of the weights learned by the GRU during fusion showed that the model assigned equal importance to all modalities when using the mean of annotations as the gold standard. However, when considering the median and EWE gold standards, the audio modality emerged as the most influential, while the text and video modalities showed comparable contributions to the overall predictions.
TABLE VIII: Time-continuous dimension prediction (CCC).
@@ VII Conclusion
This paper introduced the THERADIA WoZ corpus, an annotated multimodal corpus on affect in the context of healthcare. The corpus relies on current affective science findings from recent forms of appraisal theories, and is made available to the affective computing research community. The data consist of 39.5h of multimodal sequences from 52 healthy older participants and 9 MCI participants performing WoZ-assisted CCT. Affect inductions were performed on 32 of the healthy older participants to maximise the chance of capturing affective episodes in the data. All sequences were fully transcribed and partially annotated based on four appraisal dimensions (i.e., novelty, intrinsic pleasantness, goal conduciveness, and coping) plus the arousal dimension, as well as 23 achievement affect labels. Dimensions were annotated both in continuous time and in summarised form, while labels were only annotated in summarised form in terms of presence and intensity. Combining the results of different analyses yields a core set of ten affective labels (five positive, five negative), evidenced as highly relevant in the context of AI-assisted CCT.
Extensive baseline experiments and analyses were then conducted to model affect for both labels and dimensions using three modalities of text, audio and video. We also explored the fusion of different modalities for both computationally inexpensive feature extraction techniques and more contextual deep representations. Results showed some dependencies between the modalities and the different representations of affect. This work provides new insights into the field of affect recognition in healthcare and helps to provide additional data for benchmark comparisons.
In conclusion, we believe that this multidisciplinary work, combining recent advances in affective computing and psychology, fits perfectly with current research challenges. Indeed, we believe that such collaborations will benefit both affective computing and psychology by pushing back the boundaries of our knowledge of affect and how it can be modelled.

### v2

@@ VI-A1 Expert based representations
We used the Mel-scale Filter Banks (MFBs) [36], Term Frequency - Inverse Document Frequency (TF-IDF) [37], and Facial Action Units (FAUs) [38], as representations of the audio, textual, and visual data, respectively.
@@ VI-A2 Self-supervised representations
We used a multilingual Wav2Vec2 model that exploits both convolutional and transformer layers to distinguish positive and negative samples of speech in a contrastive approach. The model [^5] https://huggingface.co/voidful/wav2vec2-xlsr-multilingual-56 we used has a large architecture, and was trained on 56 different languages [44], including French, and then fine-tuned for ASR.
@@ VI-B Models and training strategies
To predict the time-continuous annotations of the dimensions, it was necessary to resample certain features to align with the annotation frame rate of 10 Hz. Audio representations, sampled at 100 Hz for MFBs, and 50 Hz for Wav2vec2, were averaged into 10 ms chunks to match the annotation frequency. The same strategy was applied to the FAUs and CLIP visual representations. For textual representations, we utilised BERT exclusively, as TF-IDF provides a single feature vector for the entire sequence. Textual features were aligned to the sequence length using spline interpolation, ensuring a smooth and continuous representation of the features over time. Since the Linear and MLP models did not perform well in predicting time-continuous dimensions, we explored various GRU model architectures:
@@
    One-Layer GRU: A single-layered GRU with a hidden size set to either half the input feature size or equal to the input feature size.
@@
    Three-Layers GRU: Similar to before, but with three hidden layers.
Additionally, we tested a fixed set of hidden sizes across all GRU models: 128, 256, and 512. The results indicated that a GRU model with three hidden layers, each matching the input size, performed optimally for deep representations. In contrast, the model with a hidden size of 256 produced the best results for hand-crafted representations.
For multimodal fusion, we explored both middle and late fusion techniques to address the issue of asynchronous inputs. The best performing approach for both labels and dimension summaries was decision-based fusion using an Ordinary Least Squares (OLS) regression model [^8] https://scikit-learn.org. For time-continuous dimensions, fusion was achieved through a GRU functioning as a weighted averaging layer, effectively capturing the temporal dynamics of the features. The GRU weights were optimised on the training set.
All the models were trained with the Adam optimizer and an initial learning rate of $10^{-3}$ for affective labels and $10^{-4}$ for dimensions. The batch size and gradient accumulation were set to 10 for labels and 1 for dimensions. The maximum number of training epochs was set to 50 and an early stopping strategy was used with five epochs. The models were trained using the Mean Squared Error (MSE) as the loss function, and for evaluation, the Concordance Correlation Coefficient (CCC) was calculated on all sequences [47]. All experiments were carried out using Pytorch [^9] https://pytorch.org[48], with random seeds set to zero. The computer’s operating system was Debian GNU/Linux 10, and the GPU used to train the models was an NVIDIA Quadro RTX 6000 with 23 gigabytes of memory, CUDA version 11.3.
@@ VI-C Affect prediction for the core set labels
The results for label intensity predictions, presented in Table V, reveal some strong dependencies between modalities and affective labels; see section IV.B of the supplementary material for additional details on prediction bias related to population demographics. For instance, the label “happy” was most accurately assessed via the video modality, “interested” through the textual modality, and “frustrated” and “surprised” using the audio modality with deep representations. The choice between expert-crafted and deep features also introduced notable differences in the results according to the modality. For example, the label “frustrated” achieved a significantly higher CCC (.327) with TF-IDF features compared to MFB (.151), whereas the reverse trend was observed when using SSL representations. Overall, deep representations across textual, audio, and video modalities consistently outperformed their hand-crafted counterparts on average.
The fusion of the different modalities yielded the best results on average and for most affective labels. An analysis of the fusion weights revealed that the importance of each modality varied significantly based on its representation. For instance, when using expert features, TF-IDF had a substantially greater influence on the results, with an averaged fusion coefficient of 0.8, compared to 0.1 for audio (MFBs) and video (FAUs). However, with deep representations, the contributions of audio and video surpassed that of text, with average coefficients of .28, .35, and .35 for text, audio, and video respectively.
TABLE V: Core-set label intensity prediction results (CCC)
@@
An affective label was considered present in a given sequence if at least one annotator reported an intensity value for that label. To predict these labels, we explored two modeling approaches: training a binary classification model for each affective label or deriving a threshold from the regression models. Since the binary classification approach yielded significantly poorer results compared to the thresholded-based decision from regression models, we opted for the latter in our experiments.
The results, presented in Table VI, indicate some dependencies between the best-performing modality and the affective labels, although these differences are less pronounced than in previous experiments. Overall, no significant performance gap was observed between expert and deep representations across modalities, with the exception of the audio modality. Additionally, fusion consistently enhanced performance across most scenarios.
@@ VI-D Dimensions summaries prediction
TABLE VI: Core-set label classification results (%UAR)
@@
Results are presented in table VII for two different gold standards: the average of annotations and the median; see section IV.B of the supplementary material for further details on prediction bias related to population demographics. Comparing these gold standards, the average of annotations consistently yielded the best performance across all five dimensions. In terms of individual dimensions, arousal was best predicted by the audio modality, with no significant improvement observed from fusion with the other modalities. It also had the lowest scores among the dimensions, which might be explained by the factors discussed in Section V-C. Coping was the most accurately predicted dimension, with all modalities performing similarly well and demonstrating complementary effects during fusion. Goal conduciveness was best predicted by textual representations, with fusion incorporating audio and video modalities further improving performance. Intrinsic pleasantness was most effectively captured by the video modality, with textual representations proving relevant only when computed over the entire sequence using TF-IDF. Deep audio representations also performed comparably well. Finaly, novelty was particularly challenging for the video modality, which yielded extremely low scores, while textual representations proved to be more effective in capturing this dimension.
By analysing the relative importance of each modality based on the coefficients from the linear regression model, we found that audio features contributed very little to predicting arousal, especially when using expert features. This could explain the drop in performance observed with the multimodal approach, compared to using the audio modality alone. For the others dimensions, the contributions of each modality were influenced by the complexity of the representations. When using deep features, the audio modality emerged as the most influential. However, this contribution was diminished when expert features were used, with textual information playing a more significant role.
TABLE VII: Dimension summary prediction (CCC)
@@ VI-E Continuous dimensions prediction
Results are presented in Table VIII for three gold standards: the average and the median of the time-continous annotations, and a version based on the Evaluator Weighted Estimator (EWE), which calculates a weighted mean of the annotations using the average cross-CCC between each rater’s scores as weights [49]; for further details on prediction bias based on population demographics, see section IV.B of the supplementary material. Overall, the EWE approach consistently outperformed the others two gold standards across the five dimensions.
Arousal was the best predicted dimension, but this was primarily due to the low variance observed within this dimension, as explained in the previous section. For coping, goal conduciveness and novelty, Wav2vec2 audio representations outperformed textual features, which contrasts with the findings on the summarised dimensions. This discrepancy may be attributed to the fact that textual features provide a static representation, which may not capture the continuous variations within the data. Regarding intrinsic pleasantness, the three modalities performed similarly and proved complementary during fusion, which is consistent with the previous section’s findings. Additionally, the multimodal approach consistently improved performance. An analysis of the fusion weights revealed that when using the mean of annotations as the gold standard, the model evenly distributed importance across all modalities. However, when using the median or EWE gold standards, the audio modality emerged as the most significant, while text and video contributed similarly.
An examination of the weights learned by the GRU during fusion showed that when using the mean of annotations as the gold standard, the model assigned equal importance to all modalities. In contrast, with the median and EWE gold standards, the audio modality became the most influential, while text and video modalities contributed similarly to the overall predictions.
TABLE VIII: Time-continuous dimension prediction (CCC)
@@ VII Conclusion
This paper presents the THERADIA WoZ corpus, a valuable multimodal dataset for affective computing research in healthcare. Based on current affective science and appraisal theories, the corpus includes 39.5 hours of multimodal data from 52 healthy older participants and 9 with MCI performing WoZ-assisted cognitive training. To capture affective episodes, affect inductions were applied to 32 healthy older participants. The data is fully transcribed and partially annotated across four appraisal dimensions – novelty, intrinsic pleasantness, goal conduciveness, and coping – as well as arousal, and with 23 achievement-related affect labels. These dimensions, provided in both in continuous time and summarised forms, highlights ten core affective labels of significant relevance for AI-assisted CCT applications.
In-depth baseline experiments were conducted to model affect using text, audio, and video modalities, including investigations of multimodal fusion through both computationally efficient feature and contextual deep learning representations. Results revealed key dependencies between modalities and affect representations. This work provides new insights into the field of affect recognition, offering novel insights into affect recognition in healthcare.
This work contributes to the ongoing development of affective computing in interaction – and specifically for healthcare – providing a valuable dataset for benchmark comparisons. We believe that combining the strengths of affective computing and psychology, as demonstrated in this multidisciplinary study, will significantly advance our understanding of affect and its modeling, benefiting both fields and their applications.